OpenStudy (anonymous):

Find the slope of the tangent line to the curve -4x^2-4xy-3y^3 at the point (6,2)

OpenStudy (anonymous):

same process as the last problem... how far did you get?

OpenStudy (anonymous):

so we find the derivitive of 4x^2-4xy-3y^3?

OpenStudy (anonymous):

you use implicit differentiation, yes... to find y'

OpenStudy (anonymous):

how would i set it up?

OpenStudy (anonymous):

not sure what you're asking... did you use implicit differentiation? what result did you get?

OpenStudy (anonymous):

i didn't, im confused on how to use implicit differentiation

OpenStudy (anonymous):

you treat y as some function of x. we don't really need to know what exactly that function of x actually is... we don't care in fact. It seems weird but it works. Go through term by term and differentiate with respect to x... eg.: first term (-4x^2 ) ' pretty easy: -8x second term (-4xy) ' now we have a product of x and y use the product rule: -4 ( x'y +y'x) x' is just 1 y' is y' (we don't know explicitly what function of x y is... doesn't matter, we just leave it as y') so (-4xy)' = -4(y +y'x)

OpenStudy (anonymous):

Isn't there a missing term in the equation of the curve?

OpenStudy (anonymous):

third term... use the chain rule (since y is a function of x) try it: ( -3y^3) ' =...?

OpenStudy (anonymous):

-9y^2?

OpenStudy (anonymous):

pretty good, but y' "pops out", because of the chain rule... so -9y^2 * y'

OpenStudy (anonymous):

ohh,

OpenStudy (anonymous):

anyway... if your original equation was something like -4x^2-4xy-3y^3 =C we now have the result of differentiating that equation with respect to x: -8x -4(y +y'x) -9y^2 * y' = 0

OpenStudy (anonymous):

and now we solve for y prime?

OpenStudy (anonymous):

all that's left to do is solve for y', which is just a matter of algebra :)

OpenStudy (anonymous):

yep:)

OpenStudy (anonymous):

im getting -14/3 but its coming out incorrect

OpenStudy (anonymous):

you can imagine y is some well known function of x and take the derivative and you'll see that it works... eg.: if y = sinx (-3(sin(x))^3 )' = -9(sin(x))^2 * cos(x) { -9 y ^2 * y' ) ! (I left out a 'prime' in the original, fixt)

OpenStudy (anonymous):

what did you get as the answer what i got s showing as incorrect

OpenStudy (anonymous):

nevermind, i found what i did wrong. Thanks for the help!

OpenStudy (anonymous):

cool.:)